• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kramer

    Log entry time 00:02:22 on June23,2002

    Entry number 86631

    keyword=shift summary

    1600 take over shift. Q1 just got back to field setting
    1601 run 1731 started, production
    1643 noticed roc14 'not reporting' errors, ended run 1731
    1644 run 1733 started, production
    1645 roc14 'not reporting' errors continue, ended run 1733, coda & roc14 restart
    1655 run 1734 started, production
    1725 run 1734 ends with roc14 problems, coda & roc14 restart
    1733 run 1735 started, production (magnet trip)
    1835 left q1 trip, he level faults, ramp down and reset
    1856 Q1 back, run 1736 started, production
    1904 roc14 problems end run 1736, restart
    1918 run 1738 started, production
    2015 production run 1739 started
    2114 production run 1740 started
    2222 production run 1741 started
    2238 effieciency run 1742 started (last run at this kinematics)
    2255 start setting up normalization point
    2256 right bogies not working access to power cycle 6 knife switches, after several tries, return to operational
    2340 spect set, waiting for return of beam
    2348 start run 1743 normalization run